40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2357     Entry time: Thu May 30 07:43:21 2019
Author: Anjali 
Type: Update 
Category: 2micronLasers 
Subject: Frequency noise measurement of 2 micron source using PLL 
  • The new 2 micron detector has a 50 ohm termination resistance, which is the transimpedance resistor, and hence the gain is very low 50 Ohms.
  • We observed the output of photodetector (beat note from MZI) on oscilloscope. Initially, we observed a large DC value (20 mV) and a small AC value (peak-peak=5 mV). 20 mV of DC with 50-ohm termination corresponds to a photo current of 0.4 mA.  This corresponds to optical power of 0.42 mW (responsivity = 0.95 A/W). (I'm not sure if its really as good as 0.95 A/W at 2 microns)
  • This ratio of  small AC value to large DC value says that the contrast is poor (25 %). We were then trying to improve the contrast. The RF power and RF frequency to the AOM was already set to have the maximum contrast. Prof. Rana then removed the fiber connection from laser to isolator and redid the connection. Surprisingly, the contrast became very good. We measured a AC peak-peak of 23 mV, which indicates almost 100% contrast.smiley
  • We then added a DC block and an amplifier stage (ZFL-500 LN) after the photodetector. Looking at the RMS value of the sinusoidal signal on oscilloscope, we realised that the RMS value is increasing as the thermistor resistance value set on the temperature controllerfor the laser diode increases. From the oscilloscope readout, we measured RMS value of 80 mV, 89 mV and 105 mV respectively when the thermistor resistance values are 10 kΩ, 11 kΩ and 13 kΩ. But as per the data sheet of the laser, the typical value of thermistor resistance is 10 k kΩ and the maximum value is 10.5 kΩ. Also, as per the data sheet, the thermistor temperature coefficient is -4.4 %/oC. I suppose, from the negative value, the temperature of the laser diode is reducing as the thermistor resistance value increases. Kindly correct me. This laser diode also has a current tuning coefficient of 0.01 nm/mA and temperature tuning coefficient of 0. 1 nm/ oC.
  • Attached the videos of oscilloscope signal when the thermistor resistance is 10 kΩ. From the oscilloscope trace, there is no significant amplitude fluctuation, but the frequency seems to be fluctuating a lot.
  • I have attached another video when the beat note observed on spectrum analyser (this video was captured when we had a dc block and two amplifier stage after the photo detector), but the characteristics are the same, except the power evel is different now. The peak is fluctuating a lot (sometime, the peak is even disappearing). We observed these fluctuations even with a resolution band width of 1 MHz. This indicates that the frequency of the laser diode is fluctuating by a large factor which is even greater than the maximum actuation Marconi can provide. Hence, we will not be able to lock the PLL with this laser having large frequency fluctuationfrown. We need to find another method to measure the frequency noise of the laser. One option is to perform RF herterodyne, but we don’t have a deep memory oscilloscope or ADC with large sampling rate to capture the data. Kindly give further suggestions to perform the frequency noise measurement.

RXA: we have a few options for measuring large frequency fluctuations:

  1. make an electronic delay line frequency discriminator. This is what is used at the 40m lab to track the ALS beat note. This requires the use of a medium power RF amp (ZHL-3A or similar), a splitter, 1 short cable, 1 long cable, and a mixer/LP. Two of these setups to get I & Q.
  2. a VCO with a much larger range than the Marconi - maybe 10 MHz p-p would be enough
  3. Use the heterodyne setup that Anjali mentions: a 90 deg hybrid splitter to get I & Q and then record the two mixer outputs with the Moku
Attachment 1: Oscilloscope.mp4  15.929 MB
Attachment 2: Spectrum_analyser.mp4  4.591 MB
ELOG V3.1.3-